### v2

# Simplified functional flow equation

Abstract
We adapt the precise definition of the flowing effective action in order to obtain a functional flow equation with simple properties close to physical intuition. The simplified flow equation is invariant under local gauge transformations and suitable for both euclidean and Minkowski signature and analytic continuation. The cutoff always removes fluctuations close to zeros of the inverse full propagator. A formulation of the simplified flow equation in terms of renormalized scale invariant fields permits direct access to scaling solutions and associated fixed points. Corrections to the simplified flow equation involve a field-dependent modification of the cutoff for which we discuss a systematic expansion. Truncated solutions for a scalar field theory in four dimensions suggest a new fixed point with a field-dependent coefficient of the kinetic term.
Functional flow equations for statistical systems or quantum field theories [1, 2, 3, 4] have a long history. Originally, they were derived as series of equivalent actions for which the ultraviolet cutoff is lowered continuously. This procedure integrates out fluctuations in a stepwise manner, in analogy to the block spin formalism [5]. A more recent development has shifted the focus to an average action [6] for which an infrared cutoff is introduced in the functional integral, corresponding to an effective action for averages of fields over a finite instead of infinite volume. An exact flow equation for the effective average action $\Gamma_{k}$ is based on an infrared cutoff term which is quadratic in the fields [7, 8, 9, 10, 11]. This cutoff suppresses in the functional integral all fluctuations with momenta $q^{2}$ smaller than $k^{2}$. The quadratic form of the cutoff term ensures the simple one-loop form of the functional differential equation which describes the dependence of $\Gamma_{k}$ on the “renormalization scale” $k$. Many interesting results in a wide area of fields are based on this equation [12].
Despite this success, stumbling blocks for several issues require sometimes a rather high degree of complexity of computations. The first concerns the implementation of local gauge symmetries. The exact flow equation in the background field formalism [8] involves two types of gauge fields: the background gauge field $\bar{A}_{\mu}$ and the macroscopic gauge field $A_{\mu}$. The effective action is invariant only under simultaneous gauge transformations of $\bar{A}_{\mu}$ and $A_{\mu}$. Suitable “background field identities” ensure that there is only one independent gauge field [8, 13, 14, 15]. Nevertheless, it is not straight forward to construct a gauge invariant effective action involving a single gauge field, from which correlation functions can be extracted directly by taking functional derivatives. This is particularly cumbersome for quantum gravity since both the field equations and the correlation function (“primordial fluctuations”) are derived in practice from a diffeomorphism invariant quantum effective action for a single metric field [16, 17].
A second issue concerns the analytic structure of the effective action. In principle, the flow equation can be applied both for the euclidean signature and Minkowski signature [18]. Beyond simple truncations it is often difficult to conceive an infrared cutoff that respects the analytic structure of the correlation functions [19, 20, 21, 22]. One would like to employ a cutoff which allows analytic continuation between euclidean and Minkowski signature at all stages of the flow.
Finally, the use of the exact flow equation extends to many interesting problems where large fluctuation effects arise from the neighborhood of non-vanishing “critical momenta”, as the Fermi surface for strongly correlated fermions. It is, however, not always easy to device a cutoff that first removes precisely these “critical fluctuations”, such that they are only consecutively included as the renormalization scale $k$ is lowered to zero.
In this note we explore to what extent these issues can be addressed by means of a simplification of the flow equation. The simplified flow equation aims for a setting where the effects of fluctuations are included with the presence of a type of infrared cutoff, similar to the effective average action. The corresponding flow equation should respect local gauge symmetries, be compatible with analytic continuation and realize a cutoff that automatically removes the problematic fluctuations. The prize to pay is that this flow equation may not be exact. Nevertheless, we relate this simplified flow equation to the exact flow equation of ref. [7] by choosing an appropriate cutoff and omitting certain contributions.
We propose an optimization of the definition of the flowing action for which the simplified flow equation can be considered as the starting point of a systematic expansion. We compute explicitly the general form of the corrections to the simplified flow equation. For our definition of the flowing effective action the exact flow equation can be formulated in an implicit form. The expansion around the simplified flow equation has the character of a type of loop expansion for a correction term. The simplified flow equation itself is non-perturbative. It includes already all one-loop effects exactly. Having a structure similar to the exact flow equation for the effective average action the simplified flow equation also sums up higher loops and can be used for a full non-perturbative analysis. The loop expansion only concerns corrections to the precise shape of the infrared cutoff.
Optimization by the choice of a cutoff term which is quadratic in the fluctuation field and independent of the macroscopic field has already a long tradition [23]. We aim here for a more general optimization for which the cutoff term is allowed to depend on the macroscopic fields [24, 25]. These macroscopic fields are the expectation values of the microscopic fields over which one integrates in the functional integral. The price to pay is that the one-loop form of the exact flow equation is no longer guaranteed to hold. We propose an optimized flow equation which maintains the simple one-loop form. The propagator in the loop involves, however, field-dependent corrections to the effective infrared cutoff which are, in turn, given by a loop expansion.
For a general definition of a flowing action $\Gamma_{k}$ we require that for $k=0$ all fluctuations are included, such that $\Gamma_{0}$ is the quantum effective action (or Gibbs free energy). This condition still leaves a large freedom for the choice of a functional integral that defines $\Gamma_{k}$ for $k$ different from zero. For an “optimized definition” one has to fix the optimization goals. For our purpose we ask that $\Gamma_{k}$ respects all symmetries of the microscopic physics. In case of local gauge symmetries it should be a gauge invariant functional of a single gauge field or metric. We further ask for “universality” of the cutoff in the sense that it automatically affects the “critical fluctuations”. Finally, we aim for consistency with the analytic structure. These goals do not necessarily coincide with the goal of best quantitative accuracy. As compared to the more standard effective average action the simplified flowing action may also be conceptually more distant from intuitive procedures as averaging or coarse graining. Our aim here is qualitative robustness. For other purposes the effective average action or other forms of optimization may be more appropriate.
For a single real scalar field $\varphi(x)$ the flowing action $\Gamma_{k}[\varphi]$ is a functional of this “macroscopic field”. It flows in dependence on the renormalization scale $k$. In the simplest approximation in flat Euclidean space the simplified functional flow equation reads
$k\partial_{k}\Gamma_{k}[\varphi]=k^{2n}\,\mathrm{tr}\Big{\{}\left(1-\frac{\eta}{2}\right)\left(Z^{-1}\Gamma_{k}^{(2)}[\varphi]+k^{2}\right)^{-n}\Big{\}}\ .$ (1)
For $d$ even dimensions a typical value is $n=d/2+1$. With $\Gamma_{k}^{(2)}$ the matrix of second functional derivatives of $\Gamma_{k}$ the r.h.s. of eq. (1) involves the $n$-th power of the regularized full field-dependent propagator. The quantity $Z$ denotes a suitable wave function renormalization, with anomalous dimension $\eta=-k\partial_{k}\ln Z$. Below we will generalize the wave function renormalization $Z$ to a field-dependent quantity. This replaces $Z^{-1}\Gamma_{k}^{(2)}$ by a renormalized version of the second functional derivative, which is an important characteristic of our setting.
Eq. (1) has been proposed in [26] for the Wilsonian action which differs conceptually from the flowing effective action. In this context it follows from a version of heat kernel regularization. Even though heat-kernel evaluations of the simplified flow equation are possible, we do not base our approach on a heat kernel regularization.
In order to get a first impression of the simplified flow equation one may evaluate the flow of the effective potential according to eq. (1). In momentum space one has
$\Gamma_{k}^{(2)}(q,q^{\prime})=\frac{\partial^{2}\Gamma_{k}}{\partial\varphi(-q)\partial\varphi(q^{\prime})}\ ,$ (2)
and $k^{2}$ stands for $k^{2}\delta(q,q^{\prime})$. We use conventions with $\int_{q}=\int\text{d}^{d}q(2\pi)^{-d}$, $\delta(q,q^{\prime})=(2\pi)^{d}\delta(q-q^{\prime})$, $\mathrm{tr}A(q,q^{\prime})=\int_{q,q^{\prime}}A(q,q^{\prime})\delta(q^{\prime},q)$, with $\delta^{2}(q,q^{\prime})=\Omega\delta(q,q^{\prime})$ and $\Omega$ the spacetime volume. We may take the truncation
$\Gamma_{k}^{(2)}(q,q^{\prime})=Z\big{(}q^{2}+m^{2}\big{)}\delta(q,q^{\prime})\ ,$ (3)
with $m^{2}(\varphi)$ given by the second derivative of the scalar potential $m^{2}=\partial^{2}U/\partial\varphi^{2}$. For $d=4$, $n=3$ eq. (1) yields
$k\partial_{k}\Gamma_{k}=\Omega k^{6}\left(1-\frac{\eta}{2}\right)\int_{q}\big{(}q^{2}+m^{2}+k^{2}\big{)}^{-3}\ .$ (4)
For a constant scalar field $\varphi$ one has $\Gamma_{k}=\Omega U$, with $U$ the effective potential flowing as
$k\partial_{k}U=$ $k^{6}\left(1-\frac{\eta}{2}\right)\int_{q}\big{(}q^{2}+m^{2}+k^{2}\big{)}^{-3}$
$=$ $\frac{\left(1-\frac{\eta}{2}\right)k^{6}}{32\pi^{2}\big{(}m^{2}+k^{2}\big{)}}\ .$ (5)
Neglecting $\eta$ this yields the same expression as for the standard flow equation with the Litim cutoff [23].
Simplified flow equation
We next introduce the general form of the simplified flow equation. For several fields $\varphi$ becomes a vector in field space and $\Gamma_{k}^{(2)}$ a corresponding matrix. The simplified flow equation is defined by
$k\partial_{k}\Gamma_{k}[\varphi]=k^{2n}\mathrm{tr}\Big{\{}\big{(}1-\frac{1}{2}E[\varphi]\big{)}\big{(}\Gamma_{R}^{(2)}[\varphi]+k^{2}\big{)}^{-n}\Big{\}}\ ,$ (6)
with “renormalized second functional derivative”
$\Gamma_{R}^{(2)}[\varphi]=g^{-1/2}N^{-1}[\varphi]\Gamma_{k}^{(2)}[\varphi]\big{(}N^{T}\big{)}^{-1}[\varphi]\ .$ (7)
The matrix $N[\varphi]$ generalizes $Z^{1/2}$ to a type of field-dependent wave function renormalization. Correspondingly, the matrix $E[\varphi]$ generalizes $\eta$ in eq. (1). We take
$E[\varphi]=-N^{-1}\partial_{t}\big{(}NN^{T}\big{)}\big{(}N^{T}\big{)}^{-1}[\varphi],$ (8)
such that for a $k$-dependence of $NN^{T}$ of the form $N[\varphi]N^{T}[\varphi]=A[\varphi]k^{-\eta}$ one indeed finds $E[\varphi]=\eta$. This includes the field-independent case $NN^{T}=Z$. For an arbitrary metric $g_{\mu\nu}$, with $g=\det(g_{\mu\nu})$, the factor $g^{-1/2}$ takes out the factor $g^{1/2}$ arising from $\Gamma=\int_{x}\sqrt{g}L$. In flat Euclidean space one has $g_{\mu\nu}=\delta_{\mu\nu}$ and $g=1$, while for Minkowski space $g_{\mu\nu}=\eta_{\mu\nu}$ implies $g=-1$. The factor $\sqrt{g}$ takes care of factors of $i$. The optimized flow equation (6) can therefore be used for an arbitrary signature of the metric. Analytic continuation can be performed by analytic continuation of the metric, keeping the coordinates fixed. All factors of $i$ are accounted for by the metric.
We will motivate later the definition of the simplified flow equation (6) and discuss the choice of $N[\varphi]$ for eqs. (7), (8). We also will investigate how the corrections as compared to an exact flow equation can be minimized. For the moment we first highlight a few attractive properties of eq. (6).
  1. (i)
    The infrared regulator $k^{2}$ always cuts off small values of $\Gamma_{R}^{(2)}\lesssim k^{2}$. This can concern small momenta or, in a more general context, a range of “critical momenta” for which $\Gamma_{R}^{(2)}$ is small and the fluctuation effects therefore enhanced. Infrared finiteness of the momentum integral is guaranteed for $\Gamma_{R}^{(2)}>-k^{2}$.
  2. (ii)
    Ultraviolet finiteness of the momentum integral can always be achieved by choosing large enough $n$.
  3. (iii)
    Concerning analytic continuation the only poles in the momentum integral can occur for zero eigenvalues of the matrix $\Gamma_{R}^{(2)}+k^{2}$, provided $E[\varphi]$ remains regular.
  4. (iv)
    Massive particles with $m^{2}\gg k^{2}$ decouple effectively, with their contribution suppressed by $k^{2}/m^{2}$.
In case of symmetries the matrices $N$ and $E$ are typically block diagonal for the different representations. For fermions one has an additional minus sign such that the trace becomes a supertrace. For the fermionic part one also adapts the cutoff function such that the matrix $\Gamma_{R}^{(2)}$ in eq. (6) is replaced by its square. The latter is typically quadratic in momentum, similar to $\Gamma_{R}^{(2)}$ for bosons. For applications in condensed matter physics the cutoff concerns the momenta close to the Fermi surface.
For theories with gauge symmetries an effective projection on the physical fluctuations is achieved by a “physical gauge fixing”. Contributions from gauge fluctuations and a regularized Faddeev-Popov determinant can be combined into a universal measure factor which is independent of the precise form of $\Gamma_{k}$ [24]. This measure contribution has to be added to the simplified flow equation (6). With $\Gamma_{R}^{(2)}$ involving covariant derivatives the flow equation respects the gauge symmetry, provided one employs a gauge covariant $N[\varphi]$. This points to one more central feature of the flow equation (6):
  1. (v)
    The flow equation respects all symmetries, including local gauge symmetries.
Finally, the flow equation (6) offers an important technical advantage:
  1. (vi)
    Heat kernel methods can be employed for the evaluation of the flow generator.
For a flow equation for which the flow generator (r.h.s. of eq. (6)) can be expressed as a trace over a function of the matrix $\Gamma_{R}^{(2)}$ one can employ the identity
$k\partial_{k}\Gamma_{k}=\mathrm{tr}W\big{(}\Gamma_{R}^{(2)}+k^{2}\big{)}=\int_{0}^{\infty}\text{d}z\,W(z)B(z)\ ,$ (9)
with
$B(z)=\int\frac{\text{d}s}{2\pi i}e^{sz}\mathrm{tr}\exp\Big{[}-s\big{(}\Gamma_{R}^{(2)}+k^{2}\big{)}\Big{]}\ .$ (10)
For a simple enough form of $\Gamma_{R}^{(2)}$, for example in a derivative expansion, one can use standard heat kernel expressions for the trace in eq. (10). An example is the computation of the flow of a field-dependent effective Planck mass in quantum gravity [27]. For gauge invariant $\Gamma_{k}$ the second functional derivative is gauge covariant, which makes the gauge invariance of $B(z)$ and therefore the flow generator manifest. A cutoff depending on the second functional derivative of an effective action and its relation to heat kernels has been discussed for the background field formalism in ref. [28, 29]. For other proposals of non-perturbative flow equations using heat kernels see refs. [30, 31, 32, 33, 34, 35, 26, 36].
As a second technical advantage we may state
  1. (vii)
    A suitable choice of the field-dependent wave function renormalization $N[\varphi]$ can simplify the form of $\Gamma_{R}^{(2)}$.
For many examples and a large set of truncations of $\Gamma_{k}$ this can be used to bring $\Gamma_{R}^{(2)}$ to a form involving covariant Laplacians in a simple way. One typically aims for settings where $\Gamma_{R}^{(2)}$ is block diagonal and $E$ can be approximated for each block by an anomalous dimension $\eta_{i}$. The factors $N^{-1}$ first partly diagonalize $\Gamma_{k}^{(2)}$ and subsequently perform a rescaling.
Relation to exact flow equation
The simplified flow equation is typically not an exact equation. Relating it to the exact flow equation for the effective average action [7] will produce a “correction term” to the flow equation. We will discuss in the second part of this note to which extent the precise definition of the flowing effective action can be optimized in order to minimize this correction term. Part of this optimization is the choice of the cutoff function.
Let us start for bosons with the flow equation $(\partial_{t}=k\partial_{k})$
$\partial_{t}\Gamma_{k}=\frac{1}{2}\mathrm{tr}\Big{\{}\partial_{t}R_{k}\big{(}\Gamma_{k}^{(2)}+R_{k}\big{)}^{-1}\Big{\}}\ ,$ (11)
where
$R_{k}=g^{1/2}k^{2}Nr(B)N^{T}\ ,\quad B=k^{-2}\Gamma_{R}^{(2)}\ .$ (12)
This flow equation would be exact only if $R_{k}$ is independent of $\varphi$. In general, this is not the case here since both matrices $N$ and $B$ are allowed to depend on $\varphi$. The corrections to eq. (11) for field-dependent $B$ and $N$ will be addressed below. We may consider eq. (11) as the flow equation in the background formalism for which the correction term discussed in ref. [8] has been omitted. Alternatively, it corresponds to the gauge invariant flow equation without the correction terms specified in refs. [24, 25].
We want to demonstrate that for a suitable choice of the cutoff function $R_{k}$ eq. (11) yields the simplified flow equation (6). This holds up to small corrections. We approximate
$\partial_{t}B=-2B\ ,\quad\partial_{t}r=-2B\frac{\partial r}{\partial B}\ ,$ (13)
which neglects the $k$-dependence of $\Gamma_{R}^{(2)}$ in eq. (12). In this approximation one obtains for constant $g$
$\partial_{t}R_{k}=$  $\,2g^{1/2}k^{2}N\tilde{r}N^{T}\ ,$ (14)
$\tilde{r}=$ $\,r+\frac{1}{2}N^{-1}\partial_{t}Nr+\frac{1}{2}r\partial_{t}N^{T}\big{(}N^{T}\big{)}^{-1}-B\frac{\partial r}{\partial B}\ .$
Employing
$\Gamma_{k}^{(2)}+R_{k}=g^{1/2}k^{2}N(B+r)N^{T}$ (15)
yields
$\partial_{t}\Gamma_{k}=$ $\,\mathrm{tr}\Big{\{}\Big{[}r+\frac{1}{2}N^{-1}\partial_{t}Nr+\frac{1}{2}r\partial_{t}N^{T}\big{(}N^{T}\big{)}^{-1}-B\frac{\partial r}{\partial B}\Big{]}$
×$\quad\times(B+r)^{-1}\Big{\}}$
$=$ $\,\mathrm{tr}\Big{\{}\Big{[}\left(1-\frac{1}{2}E\right)r-B\frac{\partial r}{\partial B}\Big{]}(B+r)^{-1}\Big{\}}\ .$ (16)
The second expression neglects a commutator. Together with terms neglected in eq. (13) this could be (partly) absorbed by a more elaborate definition of $E[\varphi]$.
Let us take a basis for which $B$ is block-diagonal and approximate in each block $E$ by a constant $\eta_{i}$. The contribution of a given block to the flow equation, with $\gamma_{i}=(1-\eta_{i}/2)^{-1}$, reads
$\partial_{t}\Gamma_{k}^{(i)}=\mathrm{tr}^{(i)}\Big{\{}\left(1-\frac{\eta_{i}}{2}\right)\left(r_{i}-\gamma_{i}B_{i}\frac{\partial r_{i}}{\partial B_{i}}\right)(B_{i}+r_{i})^{-1}\Big{\}}\ .$ (17)
The functions $r_{i}(x)$ are chosen such that
$(r-\gamma_{i}x\partial_{x}r)(x+r)^{-1}=(x+1)^{-n}\ .$ (18)
This yields eq. (6) with a corresponding approximation for $E$.
Our choice of the cutoff function $r$ corresponds to a solution of the differential equation (18) with the boundary condition that $r(x)$ vanishes for $x\to\infty$ and approaches a constant for $x\to 0$. We demonstrate this for $n=3$ and $\gamma=1$, where $r$ has to obey
$\frac{\text{d}r}{\text{d}x}=(1+x)^{-3}\big{[}r\big{(}3+3x+x^{2}\big{)}-1\big{]}\ .$ (19)
The solution for large $x$ is given by
$r=\frac{1}{3x^{2}}\ .$ (20)
For small $x$ one has
$r=a+(3a-1)x\ ,$ (21)
where the “integration constant” $a$ is fixed by solving eq. (19) with boundary condition (20). One may expand
$r=\frac{1}{3(1+x)^{2}}+\sum_{j=3}^{\infty}c_{j}(1+x)^{-j}\ ,$ (22)
with iterative relation for $j\geq 4$
$c_{j+1}=-\frac{1}{j+2}(c_{j}+c_{j-1})\ ,$ (23)
leading to $c_{3}=-1/12$, $c_{4}=-1/20$, $c_{5}=1/45$ etc., and
$a=\frac{1}{3}+\sum_{j=3}^{\infty}c_{j}\ .$ (24)
One can actually find an exact solution of eq. (19),
$r=1+\frac{1}{z}\left[\exp\left(-z-\frac{z^{2}}{2}\right)-1\right]\ ,\quad z=\frac{1}{1+x}\ ,$ (25)
with $a=r(x=0)=r(z=1)=\exp(-3/2)$. For $x\to-1$, $z\to\infty$, one obtains $r\to 1$ with $r+x\to(1+x)\exp\{-1/\big{(}2(1+x^{2})\big{)}\}$. The singular behavior of the pole of the flow equation for $\Gamma_{R}^{(2)}\to-k^{2}$ is related to the singular behavior of $(B+r)^{-1}$ in the flow equation (Relation to exact flow equation).
In summary, the simple analytic structure of the simplified flow equation (6) is achieved by the adaption of the cutoff function. We see no obvious reason why similar solutions of the differential equation (18) should not exist for $\gamma\neq 1$. This possibly extends to the inclusion of corrections to $\tilde{r}$. For an optimal choice of the cutoff function we believe that the dominant corrections to the simplified flow equation arises from the corrections to eq. (11) for field-dependent $R_{k}$. They add to $R_{k}$ in eq. (11) a correction term $Q_{k}$, as we will establish below in eq. (61).
Scale invariant fields and scaling form of flow
equation
We may express the flow equation in terms of renormalized scale invariant fields. In four dimensions the renormalized scale invariant scalar field $\tilde{\varphi}$ and metric $\tilde{g}_{\mu\nu}$ are given by
$\tilde{\varphi}(x)=\frac{N\varphi(x)}{k}\ ,\quad\tilde{g}_{\mu\nu}(x)=k^{2}g_{\mu\nu}(x)\ .$ (26)
We employ a scale invariant formulation with fixed coordinates, such that the scale invariant metric appears even for a flat background geometry according to
$\Gamma_{k}=$ $\int_{x}\sqrt{g(x)}\mathcal{L}(x)=k^{-4}\int_{x}\sqrt{\tilde{g}(x)}\mathcal{L}(x)$
$=$ $\int_{x}\sqrt{\tilde{g}(x)}\widetilde{\mathcal{L}}(x)\ .$ (27)
(Alternatively, one could introduce scale invariant coordinates and momenta). The flow equation at fixed scale invariant fields $\tilde{\varphi}(x)$, $\tilde{g}_{\mu\nu}(x)$ obtains from eq. (6) by a variable transformation,
$k\partial_{k}\Gamma[\tilde{\varphi},\tilde{g}_{\mu\nu}]=$ $\mathrm{tr}\Big{\{}\left(1-\frac{1}{2}\widetilde{E}\right)\left(\overline{\Gamma}^{(2)}+1\right)^{-n}\Big{\}}$ (28)
+$+\int\Big{\{}\left(1+\frac{\eta}{2}\right)\tilde{\varphi}\frac{\partial\Gamma}{\partial\tilde{\varphi}}-2\tilde{g}_{\mu\nu}\frac{\partial\Gamma}{\partial\tilde{g}_{\mu\nu}}\Big{\}}\ .$
The trace and the integral can be performed in position or momentum space, and $1$ stands for the unit matrix. The matrix $\eta(\tilde{\varphi})$ is defined by
$k\partial_{k}N(\varphi)=-\frac{1}{2}\eta(\varphi)N(\varphi)\ ,$ (29)
and we employ
$\overline{\Gamma}_{ab}^{(2)}=\tilde{g}^{-\frac{1}{2}}\frac{\partial^{2}\Gamma}{\partial\tilde{\varphi}_{a}\partial\tilde{\varphi}_{b}}\ .$ (30)
The matrices $E$ and $\widetilde{E}$ in eqs. (6) and (28) are related by
$(1-\frac{1}{2}E)(1+\Gamma_{R}^{(2)}/k^{2})^{-n}=(1-\frac{1}{2}\widetilde{E})(1+\overline{\Gamma}^{(2)})^{-n}\ .$ (31)
This absorbs possible connection terms from $\varphi$-dependent $N$ into the generalized anomalous dimension $\widetilde{E}$. For $N$ independent of $\tilde{\varphi}$ the matrices $E$ and $\widetilde{E}$ are identical. For a fixed background geometry one has $\partial\Gamma/\partial\tilde{\varphi}\sim\sqrt{\tilde{g}}$ and the dependence of $\partial^{2}\Gamma/\partial\tilde{\varphi}^{2}$ on $\sqrt{\tilde{g}}$ provides for the correct dimensional scaling. Scaling solutions obtain as solutions of the differential equation obtained by setting $\partial_{t}\Gamma_{k}[\tilde{\varphi},\tilde{g}_{\mu\nu}]=0$. There may exist different versions of scale invariant fields leading possibly to further scaling solutions.
Flow of scalar effective potential
We next demonstrate the practical use of the simplified flow equation by a computation of the flow of the effective potential of the Higgs doublet in the standard model of particle physics, and the flow of the kinetial (field-dependent coefficient of kinetic term) for a scalar field theory. For the flow of the potential we employ directly the scaling form (28) of the flow equation, which yields a very simple expression from which the flow of the quartic coupling etc. can be inferred by straightforward differentiation. In four dimensions and for $n=3$ we evaluate eq. (28) for constant $\tilde{\varphi}$ and $\tilde{g}_{\mu\nu}=k^{2}\eta_{\mu\nu}$, whereby $\Gamma=\int_{x}\sqrt{\tilde{g}}u$, with $u=U/k^{4}$ the dimensionless scalar potential. For the example of the Higgs scalar, $\tilde{\varphi}$ is replaced by a complex doublet corresponding to four real fields $\tilde{\varphi}_{\gamma}$, and we consider the invariant $\tilde{\rho}=\frac{1}{2}\sum_{\gamma}\tilde{\varphi}_{\gamma}^{2}$. One infers from eq. (28)
$k\partial_{k}u(\tilde{\rho})=\sum_{i}\frac{2-\eta_{i}}{64\pi^{2}(1+\tilde{m}_{i}^{2})}+(2+\eta)\tilde{\rho}\partial_{\tilde{\rho}}u(\tilde{\rho})-4u\ ,$ (32)
in accordance with eq. (Simplified functional flow equation).
The sum runs over all particles with renormalized dimensionless mass term $\tilde{m}_{i}^{2}=m_{i}^{2}/k^{2}$. Only particles with small enough $\tilde{m}_{i}^{2}$ contribute to the flow in a given range of $k$. For the scalar fluctuations there are three Goldstone modes with $\tilde{m}_{i}^{2}=u^{\prime}$, while the radial mode obeys $\tilde{m}^{2}=u^{\prime}+2\tilde{\rho}u^{\prime\prime}$. Here primes denote derivatives with respect to $\tilde{\rho}$. For the contribution from fermion fluctuations one has an additional minus sign and $\tilde{m}_{i}^{2}=y_{i}^{2}\tilde{\rho}$, with $y_{i}$ the Yukawa couplings. (Each Majorana fermion contributes a factor $-2$ such that for a Dirac fermion there is a factor $-4$.) For the two $W$-Bosons one finds $\tilde{m}_{i}^{2}=g_{2}^{2}\tilde{\rho}/2$ and for the $Z$-boson $\tilde{m}_{i}^{2}=g_{2}^{2}\tilde{\rho}/2+3g_{1}\tilde{\rho}/10$, with $g_{2}^{2}$ and $g_{1}^{2}$ the gauge couplings of the electroweak gauge groups $SU(2)$ and $U(1)$. For each gauge boson there is a factor three corresponding to the number of physical degrees of freedom for a massive gauge boson. The measure factor subtracts for each gauge boson the contribution of one field with $\tilde{m}_{i}^{2}=0$. At the minimum of the potential this cancels the contribution of the scalar Goldstone modes, such that only contributions from physical fluctuations remain. The contribution from physical metric fluctuations can be included in a similar way [37, 38, 39]. Eq. (Simplified functional flow equation) holds for an arbitrary constant metric, and therefore both for Euclidean or Minkowski flat space.
The flow of $\tilde{\rho}$-derivatives of $u$ obtains by differentiating both sides of eq. (32) with respect to $\tilde{\rho}$. For the quartic scalar coupling $\lambda(\tilde{\rho})=u^{\prime\prime}(\tilde{\rho})$ this yields
$k\partial_{k}$ $\lambda(\tilde{\rho})=k\partial_{k}u^{\prime\prime}(\tilde{\rho})=2\eta u^{\prime\prime}$
[…]
$-\frac{12y_{t}^{2}}{(1+y_{t}^{2}\tilde{\rho})^{3}}+\frac{3g_{2}^{4}}{2(1+g_{2}^{2}\tilde{\rho}/2)^{3}}$
$+\frac{3(g_{2}^{2}+3g_{1}^{2}/5)^{2}}{4(1+g_{2}^{2}\tilde{\rho}/2+3g_{1}^{2}\tilde{\rho}/10)^{3}}\bigg{\}}\ .$ (33)
Here we have neglected the anomalous dimensions $\eta_{i}$ in the numerators $\sim(2-\eta_{i})$, as well as $u^{\prime\prime\prime}$ and $u^{(4)}$. We only include the top quark Yukawa coupling $y_{t}$ and take $\tilde{\rho}$-independent $y_{t}$, $g_{2}$, $g_{1}$. We may evaluate the quartic scalar coupling at the potential minimum ($u^{\prime}=0$) in a range of small enough $\tilde{\rho}$, such that the mass terms in the denominators can be neglected. This approximation yields the perturbative one-loop running, provided we also employ the one-loop value for the anomalous dimension of the Higgs scalar $\eta=\big{(}6y_{t}^{2}-9g_{2}^{2}/2-9g_{1}^{2}/10\big{)}/\big{(}16\pi^{2}\big{)}$. As $k$ is lowered and $\tilde{\rho}$ increased the flow provides for appropriate thresholds for the decoupling of massive particles. Metric fluctuations are not included since they play a role only for $k$ in the vicinity of the effective Planck mass.
This rather simple computation demonstrates the power of the simplified flow equation. The effective potential is a manifestly gauge invariant quantity. In the approximation used here analytic continuation is trivial. Known one-loop results are reproduced in a simple way, supplemented by a natural description of threshold effects due to the decoupling of heavy particles. Extending the truncation in various directions proceeds by more elaborate approximations of the effective propagator. Two loop accuracy of the flow equation [40] presumably needs the correction to the simplified flow equation discussed below.
Flow of kinetic term and kinetic fixed point
As a second practical demonstration we compute the flow of the kinetial for the quantum field theory of a single real scalar field $\varphi$. The kinetial $K(\varphi)$ is the field-dependent coefficient of the term with two derivatives. This field-dependence induces interactions which can lead to complex features if $K$ is not close to a constant [41]. For $K\sim\varphi^{-2}$ the model exhibits scale symmetry with respect to multiplicative rescaling of $\varphi$. This symmetry requires a flat potential. One would like to implement a flow equation respecting this symmetry in order to investigate the possible existence of a fixed point exhibiting this scale symmetry. A scale-invariant cutoff is necessarily field-dependent. The simplified flow equation offers a tool for this investigation.
Let us consider a flowing action of the form
$\Gamma_{k}=\int_{x}\sqrt{g}\left\{\frac{1}{2}K(\varphi)\partial^{\mu}\varphi\partial_{\mu}\varphi+U(\varphi)\right\}\ .$ (34)
The second functional derivative reads
$g^{-1/2}\Gamma_{k}^{(2)}=$ $\,-\Big{[}K\partial^{\mu}\partial_{\mu}+\partial_{\varphi}K\partial^{\mu}\varphi\partial_{\mu}+\frac{1}{2}\partial_{\varphi}^{2}K\partial^{\mu}\varphi\partial_{\mu}\varphi$
$\,+\partial_{\varphi}K\partial^{\mu}\partial_{\mu}\varphi\Big{]}+\partial_{\varphi}^{2}U\ .$ (35)
We define $\Gamma_{R}^{(2)}$ by
$\Gamma_{k}^{(2)}=g^{1/2}K^{1/2}\Gamma_{R}^{(2)}K^{1/2}\ ,$ (36)
where
$\Gamma_{R}^{(2)}=-\partial^{\mu}\partial_{\mu}-C+K^{-1}\partial_{\varphi}^{2}U\ ,$ (37)
with
$C=\frac{1}{4}\big{(}\partial_{\varphi}\ln K\big{)}^{2}\partial^{\mu}\varphi\partial_{\mu}\varphi+\frac{1}{2}\partial_{\varphi}\ln K\partial^{\mu}\partial_{\mu}\varphi\ .$ (38)
One identifies in eq. (7) $N=K^{1/2}$ and for eq. (8) $E=-\partial_{t}\ln K$. For a non-trivial field dependence of the kinetial the field-dependent renormalization factor $N(\varphi)$ in the simplified flow equation (6), (7) plays an important role.
We next evaluate the flow equation for $K(\varphi)$ from a derivative expansion of the flowing action, using simple heat kernel methods. For the heat kernel expansion we employ
$\mathrm{tr}\Big{\{}\exp\big{\{}s\big{(}g^{\mu\nu}\partial_{\mu}\partial_{\nu}-m^{2}(x)\big{)}\big{\}}\Big{\}}$
$=$ $\,\frac{1}{16\pi^{2}}\int_{x}\sqrt{g}\exp\big{(}-sm^{2}(x)\big{)}$
×$\quad\times\big{(}s^{-2}+c_{1}\partial^{\mu}\partial_{\mu}m^{2}+c_{2}s\partial^{\mu}m^{2}\partial_{\mu}m^{2}+\dots\big{)}$ (39)
with
$c_{1}=-\frac{1}{6}\ ,\quad c_{2}=\frac{1}{12}\ .$ (40)
The dots indicate terms with more than two derivatives or derivatives acting on the metric. For $n=3$ eq. (9) yields
$\partial_{t}\Gamma_{k}=$ $\,\frac{k^{6}}{16\pi^{2}}\left(1-\frac{\eta}{2}\right)\int_{x}\sqrt{g}\int_{0}^{\infty}\text{d}z\,z^{-3}$
×$\times\int\frac{\text{d}s}{2\pi i}$ $\,e^{s(z-k^{2}-m^{2}(x))}\big{(}s^{-2}+c_{1}\partial^{\mu}\partial_{\mu}m^{2}+c_{2}s\partial^{\mu}m^{2}\partial_{\mu}m^{2}\big{)}$
$=$ $\,\frac{1-\eta/2}{16\pi^{2}}\int_{x}\sqrt{g}\bigg{(}\frac{k^{6}}{2(k^{2}+m^{2})}+\frac{c_{1}\partial^{\mu}\partial_{\mu}m^{2}k^{6}}{(k^{2}+m^{2})^{3}}$
$\,+\frac{3c_{2}\partial^{\mu}m^{2}\partial_{\mu}m^{2}k^{6}}{(k^{2}+m^{2})^{4}}\bigg{)}\ .$ (41)
For homogeneous $m^{2}$ we recover the flow of the potential in eq. (Simplified functional flow equation).
With $m^{2}=\partial_{\varphi}^{2}U/K-C$ one employs for the second order of the derivative expansion
$\partial_{\mu}m^{2}=\big{(}\partial_{\varphi}^{3}UK^{-1}-\partial_{\varphi}^{2}U\partial_{\varphi}KK^{-2}\big{)}\partial_{\mu}\varphi\ .$ (42)
Writing eq. (Flow of kinetic term and kinetic fixed point) in the form
$\partial_{t}\Gamma_{k}=\int_{x}\sqrt{g}\Big{\{}\partial_{t}U+\frac{1}{2}\partial_{t}K\partial^{\mu}\varphi\partial_{\mu}\varphi\Big{\}}\ ,$ (43)
one obtains the flow of the potential
$\partial_{t}U=\frac{(1-\eta/2)k^{6}}{32\pi^{2}(k^{2}+\partial_{\varphi}^{2}UK^{-1})}\ .$ (44)
The flow equation for the kinetial is extracted by partial integration as
$\partial_{t}K=$ $\,\frac{(1-\eta/2)k^{6}}{16\pi^{2}}\bigg{[}\frac{(\partial_{\varphi}\ln K)^{2}-2\partial_{\varphi}^{2}\ln K}{4(k^{2}+\partial_{\varphi}^{2}UK^{-1})^{2}}$
$\,+\frac{(\partial_{\varphi}^{3}UK^{-1}-\partial_{\varphi}^{2}U\partial_{\varphi}KK^{-2})\partial_{\varphi}\ln K}{(k^{2}+\partial_{\varphi}^{2}UK^{-1})^{3}}$
$\,-\frac{(\partial_{\varphi}^{3}UK^{-1}-\partial_{\varphi}^{2}U\partial_{\varphi}KK^{-2})^{2}}{2(k^{2}+\partial_{\varphi}^{2}UK^{-1})^{4}}\bigg{]}\ .$ (45)
The simplified flow equation permits a straightforward computation of the flow of both the potential and the kinetial for an arbitrary field-dependent kinetial $K(\varphi)$.
We are interested in possible scaling solutions of the system (44), (Flow of kinetic term and kinetic fixed point) which can be associated to fixed points. It is striking that the flow equation (43)- (Flow of kinetic term and kinetic fixed point) admits a scaling solution with flat potential and kinetial $\sim\varphi^{-2}$
$U=u_{0}k^{4}\ ,\quad K=\frac{\kappa_{0}}{\varphi^{2}}\ ,\quad u_{0}=\frac{1}{128\pi^{2}}\ .$ (46)
This fixed point realizes the exact global scale symmetry $\varphi\to\beta\varphi$. Indeed, the simplified flow equation (6) respects this symmetry. With $\Gamma_{k}^{(2)}$ scaling $\sim\beta^{-2}$ this holds provided $N$ scales $\sim\beta^{-1}$, such that $\Gamma_{R}^{(2)}$ and $E$ are invariant. The field-dependence of the cutoff $\mathcal{R}_{k}\sim K\sim\varphi^{-2}$ in eq. (11) is crucial for the compatibility with this scale symmetry. This fixed point may be called “kinetic fixed point”. As a second fixed point one has the trivial fixed point for which one has in eq. (46) a constant kinetial $K=K_{0}$. This second fixed point is invariant under a constant shift in $\varphi$.
It will be interesting to find out if the kinetic fixed point with multiplicative scale symmetry (46) survives an extended truncation and the inclusion of corrections to the simplified flow equation. If this is the case, it offers the interesting prospect of a non-trivial UV-complete scalar field theory in four dimensions. One expects the existence of crossover trajectories between the scale invariant and the trivial fixed point. For models on the crossover trajectory one expects the presence of interactions. For a crossover trajectory any point on the trajectory can be realized as a UV-complete theory. If a fixed point with the properties (46) can be established, this may overcome the old “triviality problem” for scalar theories in four dimensions.
Exact flow equation for field-dependent cutoff
For a cutoff function $R_{k}$ which depends on the macroscopic field $\varphi$ the flow equation (11) is no longer exact. There are correction terms ${}\sim\partial_{\varphi}R_{k}$ that we will derive next. In a loop expansion they are of higher order. An optimization of the definition of the flowing action minimizes the importance of the correction term. For this purpose we define the flowing action $\Gamma_{k}[\varphi]$ by a functional identity
$\Gamma_{k}[\varphi]=-\ln\int\mathcal{D}\chi\,\exp\big{\{}-\bar{S}_{k}[\chi,\varphi]\big{\}}\ ,$ (47)
where $\bar{S}[\chi,\varphi]$ is related to the classical action $S[\chi]$ by the addition of a cutoff and a source term,
$\bar{S}_{k}[\chi,\varphi]=$ $\,S[\chi]+\frac{1}{2}(\chi-\varphi)R_{k}(\varphi)(\chi-\varphi)$
$-\left(\frac{\partial\Gamma_{k}}{\partial\varphi}+L_{k}(\varphi)\right)(\chi-\varphi)\ .$ (48)
We employ here a notation where $\varphi$ stands for a vector with components $\varphi_{\alpha}$, with $\alpha$ labeling both position or momentum and internal indices. The cutoff $R_{k}$ is a matrix with elements $(R_{k})_{\alpha\beta}$, and products of vectors stand for scalar products, $\chi\varphi=\chi_{\alpha}\varphi_{\alpha}$. For $k\to 0$ both $R_{k}$ and $L_{k}$ vanish and $\Gamma_{k\to 0}$ equals the quantum effective action, as usually defined by a Legendre transform of the Schwinger functional. For $\partial R_{k}/\partial\varphi=0$ one has $L_{k}=0$ and our definition of the flowing action equals the effective average action [7].
We define $L_{k}(\varphi)$ such that the macroscopic field $\varphi$ equals the expectation value of the microscopic field
$\langle\chi\rangle=\varphi\ .$ (49)
Here expectation values are formed with the weight factor $\exp\big{(}-\bar{S}_{k}\big{)}$,
$\langle A\rangle=\frac{\int\mathcal{D}\chi\,Ae^{-\bar{S}_{k}}}{\int\mathcal{D}\,e^{-\bar{S}_{k}}}\ .$ (50)
The logarithmic $k$-derivative of eq. (47) at fixed $\varphi$ yields the exact flow equation
$\partial_{t}\Gamma_{k}=\frac{1}{2}\mathrm{tr}\big{\{}\partial_{t}R_{k}G\big{\}}\ ,$ (51)
which involves the connected two-point function
$G_{\alpha\beta}=\langle(\chi_{\alpha}-\varphi_{\alpha})(\chi_{\beta}-\varphi_{\beta})\rangle\ .$ (52)
We need an expression for $L_{k}(\varphi)$ and a relation between $G$ and functional derivatives of $\Gamma_{k}$. Taking a derivative of eq. (47) with respect to $\varphi_{\alpha}$ yields
$\frac{\partial\Gamma_{k}}{\partial\varphi_{\alpha}}=$ $\,\left\langle\frac{\partial\bar{S}_{k}}{\partial\varphi_{\alpha}}\right\rangle=\frac{\partial\Gamma_{k}}{\partial\varphi_{\alpha}}-M_{\alpha\beta}(\langle\chi_{\beta}\rangle-\varphi_{\beta})$
$+L_{k,\alpha}+\frac{1}{2}\mathrm{tr}\left\{\frac{\partial R_{k}}{\partial\varphi_{\alpha}}G\right\}\ ,$ (53)
with
$M_{\alpha\beta}=\big{(}\Gamma_{k}^{(2)}+R_{k}\big{)}_{\alpha\beta}+\frac{\partial L_{k,\beta}}{\partial\varphi_{\alpha}}\ .$ (54)
The relation (49) is obeyed for the choice
$L_{k,\alpha}=-\frac{1}{2}\mathrm{tr}\left\{\frac{\partial R_{k}}{\partial\varphi_{\alpha}}G\right\}\ .$ (55)
Taking further a $\varphi$-derivative of eq. (49) one finds
$\delta_{\alpha\beta}=$ $\,\frac{\partial\langle\chi_{\beta}\rangle}{\partial\varphi_{\alpha}}=-\left\langle\frac{\partial\bar{S}_{k}}{\partial\varphi_{\alpha}}(\chi_{\beta}-\varphi_{\beta})\right\rangle$
$=$ $\,M_{\alpha\gamma}G_{\gamma\beta}-\frac{1}{2}\frac{\partial R_{k,\gamma\delta}}{\partial\varphi_{\alpha}}H_{\gamma\delta\beta}\ ,$ (56)
which involves the connected three-point function
$H_{\gamma\delta\beta}=\big{\langle}(\chi_{\gamma}-\varphi_{\gamma})(\chi_{\delta}-\varphi_{\delta})(\chi_{\beta}-\varphi_{\beta})\big{\rangle}\ .$ (57)
One infers the identity
$G=\big{(}\Gamma_{k}^{(2)}+R_{k}+Q_{k}\big{)}^{-1}\ ,$ (58)
with
$Q_{k,\alpha\beta}=-\frac{1}{2}\left[\frac{\partial}{\partial\varphi_{\alpha}}\mathrm{tr}\left\{\frac{\partial R_{k}}{\partial\varphi_{\beta}}G\right\}+\mathrm{tr}\left\{\frac{\partial R_{k}}{\partial\varphi_{\alpha}}\tilde{H}_{\beta}\right\}\right]\ .$ (59)
Here the matrix $\tilde{H}_{\beta}$ has elements
$\tilde{H}_{\beta,\gamma\delta}=H_{\gamma\delta\varepsilon}G^{-1}_{\varepsilon\beta}\ .$ (60)
Since the matrices $G$, $\Gamma_{k}^{(2)}$ and $R_{k}$ are symmetric, we symmetrize $Q_{k}$.
We arrive at the final form of the exact flow equation
$\partial_{t}\Gamma_{k}=\frac{1}{2}\mathrm{tr}\Big{\{}\partial_{t}R_{k}\big{(}\Gamma_{k}^{(2)}+R_{k}+Q_{k}\big{)}^{-1}\Big{\}}\ .$ (61)
The corrections to the flow equation (11) appear in the form of a field-dependent correction $Q_{k}$ to the infrared cutoff $R_{k}$ in the effective $k$- and field-dependent propagator. As long as $Q_{k}$ remains small as compared to $\Gamma_{k}^{(2)}+R_{k}$ one only finds small corrections. In order to establish the reliability of the simplified flow equation we have to compute the size of $Q_{k}$. This is done in the following. While somewhat technical, we consider this part as an important progress as compared to earlier papers on the gauge invariant flow equation where only qualitative estimates have been given.
Corrections to the simplified flow equation
In order to close the exact flow equation (61) we need to express $Q_{k}$ in terms of $\Gamma_{k}$ and $R_{k}$. This closure needs a relation between $\tilde{H}_{\beta}$ and $G$ or $Q$ and an expression for $\partial\tilde{G}/\partial\varphi_{\alpha}$. One may take a $\varphi$-derivative of the relation,
$G_{\gamma\delta}\big{(}\Gamma_{k}^{(2)}+R_{k}+Q_{k}\big{)}_{\delta\varepsilon}=\delta_{\gamma\varepsilon}\ ,$ (62)
which expresses $\partial_{\alpha}G_{\gamma\delta}$ in terms of the third functional derivative of $\Gamma_{k}$,
$\partial_{\alpha}G_{\gamma\delta}=$ $\,-G_{\gamma\varepsilon}\partial_{\alpha}\big{(}\Gamma_{k}^{(2)}+R_{k}+Q_{k}\big{)}_{\varepsilon\eta}G_{\eta\delta}\ ,$
$\partial_{\alpha}\Gamma^{(2)}_{k,\varepsilon\eta}=$ $\,\partial_{\alpha}\partial_{\varepsilon}\partial_{\eta}\Gamma_{k}=\big{(}\Gamma^{(2)}_{k}\big{)}_{\alpha\varepsilon\eta}\ .$ (63)
(We use here and from now on a notation $\partial_{\alpha}F=\partial F/\partial\varphi_{\alpha}$.) On the other hand, we may take a $\varphi$-derivative of the defining relation (52),
$\partial_{\alpha}G_{\gamma\delta}=$ $\,-\big{\langle}\partial_{\alpha}\bar{S}_{k}\big{[}(\chi_{\gamma}-\varphi_{\gamma})(\chi_{\delta}-\varphi_{\delta})-G_{\gamma\delta}\big{]}\big{\rangle}$ (64)
$=$ $\,M_{\alpha\eta}H_{\eta\gamma\delta}-\frac{1}{2}\partial_{\alpha}R_{k,\varepsilon\eta}\big{(}V_{\varepsilon\eta\gamma\delta}-G_{\varepsilon\eta}G_{\gamma\delta}\big{)}\ ,$
which contains the four-point function
$V_{\varepsilon\eta\gamma\delta}=\big{\langle}(\chi_{\varepsilon}-\varphi_{\varepsilon})(\chi_{\eta}-\varphi_{\eta})(\chi_{\gamma}-\varphi_{\gamma})(\chi_{\delta}-\varphi_{\delta})\big{\rangle}\ .$ (65)
Writing $\partial_{\alpha}G_{\gamma\delta}$ in the form
$\partial_{\alpha}G_{\gamma\delta}=$ $\,\big{(}\tilde{H}_{\alpha}\big{)}_{\gamma\delta}+N_{\alpha\gamma\delta}\ ,$
$N_{\alpha\gamma\delta}=$ $\,\frac{1}{2}\mathrm{tr}\big{\{}\partial_{\alpha}R_{k}\tilde{H}_{\eta}\big{\}}H_{\eta\gamma\delta}-\frac{1}{2}\mathrm{tr}\big{\{}\partial_{\alpha}R_{k}\tilde{V}_{\gamma\delta}\big{\}}$
$+\frac{1}{2}\mathrm{tr}\big{\{}\partial_{\alpha}R_{k}G\big{\}}G_{\gamma\delta}\ ,$
$\big{(}\tilde{V}_{\gamma\delta}\big{)}_{\varepsilon\eta}=$ $\,V_{\varepsilon\eta\gamma\delta}\ ,$ (66)
one infers for $Q_{k}$ the expression
$\big{(}Q_{k}\big{)}_{\alpha\beta}=$ $\,-\frac{1}{2}\mathrm{tr}\big{\{}\partial_{\alpha}\partial_{\beta}R_{k}G+\partial_{\alpha}R_{k}\partial_{\beta}G+\partial_{\beta}R_{k}\partial_{\alpha}G\big{\}}$
$+D_{k,\alpha\beta}\ ,$ (67)
with
$D_{k,\alpha\beta}=\frac{1}{2}\partial_{\alpha}R_{k,\gamma\delta}N_{\beta\gamma\delta}\ .$ (68)
This procedure does not close since $D_{k}$ involves the unknown four-point function $V$ through eq. (Corrections to the simplified flow equation). One could continue by taking higher $\varphi$-derivatives.
We may truncate $D_{k}=0$ and consider the resulting form of $Q_{k}$ in eq. (Corrections to the simplified flow equation) as a first term in some expansion. Higher orders in this expansion take higher $\varphi$-derivatives and include the effects of higher correlation functions. This expansion leads to a type of loop expansion for $Q_{k}$. Each trace amounts to a momentum integral in a loop. We conclude that $Q_{k}$ is of one-loop order, while $D_{k}$ involves two loops since $N_{\beta\gamma\delta}$ is already of one-loop order. Each loop order implies an additional factor $\partial_{\alpha}R_{k}$. We observe that $Q_{k}$ is loop-suppressed as compared to $\Gamma_{k}^{(2)}+R_{k}$.
In lowest order of an iterative computation of $Q_{k}$ we can neglect $D_{k}$ and approximate in eq. (Corrections to the simplified flow equation) $G$ by $\big{(}\Gamma_{k}^{(2)}+R_{k}\big{)}^{-1}$,
$Q_{k}=$ $\,Q_{k}^{(1)}+Q_{k}^{(2)}\ ,$
$Q^{(1)}_{k,\alpha\beta}=$ $\,-\frac{1}{2}\mathrm{tr}\Big{\{}\partial_{\alpha}\partial_{\beta}R_{k}\big{(}\Gamma_{k}^{(2)}+R_{k}\big{)}^{-1}\Big{\}}\ ,$ (69)
$Q^{(2)}_{k,\alpha\beta}=$ $\,-\frac{1}{2}\mathrm{tr}\Big{\{}\partial_{\alpha}R_{k}\partial_{\beta}\big{(}\Gamma_{k}^{(2)}+R_{k}\big{)}^{-1}+\alpha\leftrightarrow\beta\Big{\}}\ .$
The correction $Q^{(2)}_{k}$ has a term proportional to the third derivative of $\Gamma_{k}$, $\partial_{\beta}\big{(}\Gamma_{k}^{(2)}\big{)}_{\gamma\delta}=\Gamma^{(3)}_{k,\beta\gamma\delta}$.
The correction $Q_{k}$ in the exact flow equation (61) fits well into the structure of the simplified flow equation. One simply replaces $\Gamma_{k}^{(2)}$ by $\Gamma_{k}^{(2)}+Q_{k}$. Correspondingly, we choose the cutoff function such that $r$ is now a function of $x$
$x=k^{-2}g^{-1/2}N^{-1}\big{(}\Gamma_{k}^{(2)}+Q_{k}\big{)}\big{(}N^{T}\big{)}^{-1}\ .$ (70)
This replaces in the flow equation (6) the matrix $\Gamma_{R}^{(2)}$ by $k^{2}x=\Gamma_{R}^{(2)}+g^{-1/2}N^{-1}Q_{k}\big{(}N^{T}\big{)}^{-1}$. Including the correction the flow equation (6) is extended to
$\partial_{t}\Gamma_{k}=\mathrm{tr}\Big{\{}\big{(}1-\frac{1}{2}E\big{)}\big{(}B+x_{Q}+1\big{)}^{-n}\Big{\}}\ ,$ (71)
where
$x_{Q}=g^{-1/2}k^{-2}N^{-1}Q_{k}\big{(}N^{T}\big{)}^{-1}\ .$ (72)
As long as $x_{Q}$ remains small as compared to $1+B$ the leading order approximation (6) is valid. In the limit of constant $x_{Q}$ the analytic structure of the simplified flow equation only changes by a shift of the poles to $B=-(1+x_{Q})$. This may be seen as an “additive renormalization” of $k^{2}$ in the regularized propagator.
Lowest order correction
In lowest order the correction $x_{Q}$ inserts in eq. (72) the approximation (Corrections to the simplified flow equation). For an estimate of the size of $x_{Q}$ we neglect first $\partial_{\alpha}N$, $\partial_{\alpha}g$ as well as commutators of $x$ and $\partial_{\alpha}x$, such that
$\partial_{\alpha}\partial_{\beta}R_{k}=g^{1/2}k^{2}N\big{(}\partial_{\alpha}\partial_{\beta}xr^{\prime}+\partial_{\alpha}x\partial_{\beta}xr^{\prime\prime}\big{)}N^{T}\ ,$ (73)
where primes denote here derivatives with respect to $x$. In leading order we can replace $\partial_{\beta}x$ by $\partial_{\beta}B$ and obtain
$Q^{(1)}_{k,\alpha\beta}=-\frac{1}{2}\mathrm{tr}\Big{\{}\big{(}\partial_{\alpha}\partial_{\beta}Br^{\prime}+\partial_{\alpha}B\partial_{\beta}Br^{\prime\prime}\big{)}\big{(}B+r\big{)}^{-1}\Big{\}}\ .$ (74)
The expression (74) can be related to the renormalized dimensionless 1PI-three- and four-point functions
$\Gamma^{(3)}_{R,\alpha\gamma\delta}=$ $\,k^{2}N^{-1}_{\alpha\alpha^{\prime}}\partial_{\alpha^{\prime}}B_{\gamma\delta}\ ,$
$\Gamma^{(4)}_{R,\alpha\beta\gamma\delta}=$ $\,k^{2}N^{-1}_{\alpha\alpha^{\prime}}N^{-1}_{\beta\beta^{\prime}}\partial_{\alpha^{\prime}}\partial_{\beta^{\prime}}B_{\gamma\delta}\ ,$ (75)
implying
$x^{(1)}_{Q,\alpha\beta}=$ $\,-\frac{1}{2}g^{-1/2}\bigg{\{}k^{-4}\Gamma^{(4)}_{R,\alpha\beta\gamma\delta}\big{[}r^{\prime}(B+r)^{-1}\big{]}_{\delta\gamma}$
$+k^{-6}\Gamma^{(3)}_{R,\alpha\gamma\varepsilon}\Gamma^{(3)}_{R,\beta\varepsilon\delta}\big{[}r^{\prime\prime}(B+r)^{-1}\big{]}_{\delta\gamma}\bigg{\}}\ .$ (76)
The momentum integrals of the loop corresponding to the trace (sum over $\gamma$ ,$\delta$) are both ultraviolet and infrared finite since $r^{\prime}$ or $r^{\prime\prime}$ decay rapidly for large $x$, and $r$ approaches a constant for $x\to 0$. For homogeneous macroscopic fields $x_{Q}$ is diagonal in momentum space, similar to $r$. As compared to $r$ it is suppressed by a loop factor $(16\pi^{2})^{-1}$ and it involves the renormalized couplings (Lowest order correction). We conclude that unless the renormalized couplings (Lowest order correction) are very large the corrections to the simplified flow equation (6) are small and the iterative loop expansion seems to converge. The situation for the contribution from $Q_{k}^{(2)}$ is similar.
As an example we consider a single scalar field. For a homogeneous macroscopic field $\varphi$ one has in momentum space
$\Gamma_{R}^{(4)}(p_{1},p_{2},p_{3},p_{4})=$  $\,3\lambda_{R}\delta(p_{1}-p_{2}+p_{3}-p_{4})\ ,$
$\Gamma_{R}^{(2)}(p_{1},p_{2})=$ $\,\big{(}p_{1}^{2}+m_{R}^{2}\big{)}\delta(p_{1}-p_{2})\ .$ (77)
Both $m_{R}^{2}$ and $\lambda_{R}$ may depend on $\varphi$, while we assume here for simplicity that they do not depend on momentum. We define
$\big{[}r^{\prime}(B)\big{(}B+r(B)\big{)}^{-1}\big{]}(p_{1},p_{2})$
$=$ $\,f(B)(p_{1},p_{2})=f\left(\frac{p_{1}^{2}+m_{R}^{2}}{k^{2}}\right)\delta(p_{1}-p_{2})\ ,$ (78)
and obtain for the first term in eq. (Lowest order correction)
$\tilde{x}^{(1)}_{Q}(p_{3},p_{4})=$ $\,-\frac{3}{2}g^{-1/2}k^{-4}\lambda_{R}\delta(p_{3}-p_{4})\int_{q}f\left(\frac{q^{2}+m_{R}^{2}}{k^{2}}\right)$
$=$ $\,-\frac{3\lambda_{R}}{16\pi^{2}}h\left(\frac{m_{R}^{2}}{k^{2}}\right)\delta(p_{3}-p_{4})\ .$ (79)
With $\tilde{m}^{2}=m_{R}^{2}/k^{2}$ one has
$h(\tilde{m}^{2})=\int_{\tilde{m}^{2}}^{\infty}\text{d}x^{\prime}\,(x^{\prime}-\tilde{m}^{2})\partial_{x}r(x^{\prime})\big{(}x^{\prime}+r(x^{\prime})\big{)}^{-1}\ .$ (80)
The function $h(\tilde{m}^{2})$ becomes a constant for $\tilde{m}^{2}=0$, and decreases rapidly for $\tilde{m}^{2}\gg 1$ due to the fast decay of $\partial_{x}r$ in eq. (19). We observe the expected suppression factor $\lambda_{R}/(16\pi^{2})$. As compared to $B+1\approx q^{2}/k^{2}+\tilde{m}^{2}+1$ the correction $\tilde{x}^{(1)}_{Q}$ becomes tiny for large $\tilde{m}^{2}$ and affects the momentum integral (71) only for $q^{2}\approx k^{2}$. Similar considerations hold for the other parts of the correction term.
We conclude that the correction term has only a small influence on the flow equation for the effective potential as long as $\lambda_{R}/(16\pi^{2})$ remains small. For quantitative precision it cannot be neglected, however. Besides the small factor $\lambda_{R}/(16\pi^{2})$ the relative contribution to the $\varphi$-dependence of the potential is suppressed by $\partial h/\partial\tilde{m}^{2}$.
Robustness of kinetic fixed point
Let us next ask how corrections affect the possible kinetic fixed point (46). If the classical action is invariant under the constant scaling $\chi\to\beta\chi$, we can choose a cutoff $R_{k}$ which scales $\sim\beta^{-2}$, e.g. $\sim\varphi^{-2}$. If the flowing action $\Gamma_{k}$ is invariant under this rescaling up to a field-independent constant, one has $\partial\Gamma_{k}/\partial\varphi\sim\beta^{-1}$ and $L_{k}\sim\beta^{-1}$. One concludes that $\bar{S}_{k}$ in eq. (Exact flow equation for field-dependent cutoff) is scale invariant, and therefore $\Gamma_{k}$ is scale invariant up to a field- and $k$-independent constant which arises from the measure $\int\mathcal{D}\chi$ in eq. (47). Thus the exact flow equation (51), (61) is scale invariant. One may verify that the correction term does not perturb scale invariance since $Q_{k}$ scales $\sim\beta^{-2}$.
From the scale invariance of the exact flow equation for scale invariant $\Gamma_{k}$ one infers that the manifold of scale-invariant flowing actions constitutes a partial fixed point (or better fixed surface). Starting at a point on this surface all flow trajectories remain within this surface. For the investigation of a scale-invariant fixed point a restriction to this surface is therefore sufficient. This implies immediately a flat potential and a kinetial $K=\kappa(k)\varphi^{-2}$. An exact fixed point requires that the flow of $\kappa(k)$ has a fixed point.
With respect to the global scaling $\varphi\to\beta\varphi$ with invariant $g_{\mu\nu}$ an invariant flowing action takes in flat space the general form
$\Gamma_{k}=\int_{x}\sqrt{g}\left\{\frac{\kappa(k)}{2\varphi^{2}}\partial^{\mu}\varphi\partial_{\mu}\varphi+U_{0}(k)+\dots\right\}\ ,$ (81)
where dots denote terms with four or more derivatives of $\ln(\varphi)$. With $N[\varphi]=\sqrt{\kappa}/|\varphi|$, $E[\varphi]=-\partial_{t}\ln\kappa=\eta$, one finds that $\Gamma_{R}^{(2)}$ is independent of $\kappa$ and $k$,
$\Gamma_{R}^{(2)}=-\partial^{\mu}\partial_{\mu}+\partial^{\mu}\partial_{\mu}\ln\varphi\ ,$ (82)
such that eq. (13) holds. In lowest order the simplified flow equation yields with with eq. (Flow of kinetic term and kinetic fixed point) a vanishing flow $\partial_{t}\kappa=0$, implying a line of fixed points (46) with free parameter $\kappa_{0}$.
The correction term adds to $\Gamma_{R}^{(2)}$ in eq. (37) a contribution
$\tilde{Q}_{k}=$ $\,g^{-1/2}\kappa^{-1}\varphi Q_{k}\varphi$
$=$ $\,\frac{k^{2}}{16\pi^{2}\kappa}\big{(}f_{1}k^{2}+f_{2}\partial^{\mu}\ln\varphi\partial_{\mu}+f_{3}\partial^{\mu}\ln\varphi\partial_{\mu}\ln\varphi$
$\quad\quad\quad+f_{4}\partial^{\mu}\partial_{\mu}\ln\varphi\big{)}\ ,$ (83)
with field-independent dimensionless functions $f_{j}(-\partial^{\mu}\partial_{\mu}/k^{2})$, and $x_{Q}=\tilde{Q}/k^{2}$. Here the general form of the second expression uses scale invariance of $\tilde{Q}_{k}$ and truncates in second order of a derivative expansion. As an example, we take $f_{2}=0$ and constant $f_{1}$, $f_{3}$, $f_{4}$. Then the terms in eq. (Flow of kinetic term and kinetic fixed point) which involve $\partial_{\mu}\tilde{Q}_{k}$ contain more than two derivatives and do not contribute to the flow of $K$. The kinetial flows only for $f_{3}\neq 0$,
$\partial_{t}\kappa=-\frac{f_{3}k^{4}}{(16\pi^{2})^{2}\kappa}\left(1-\frac{\eta}{2}\right)\left(1+\frac{f_{1}k^{2}}{16\pi^{2}\kappa}\right)^{-2}\ .$ (84)
Counting factors of $\kappa$ one finds that $Q_{k}$ has no factor $\kappa$ and therefore the functions $f_{j}$ are independent of $\kappa$. The coefficients $f_{j}$ can be computed in lowest order from eq. (Corrections to the simplified flow equation). With $N=\sqrt{\kappa}/\varphi$ the terms $\sim\partial_{\alpha}N$ can no longer be neglected. Independently of the precise values of $f_{j}$ the flow of $\kappa$ becomes very slow for large $\kappa/k^{2}$. There is an exact fixed point at $\kappa^{-1}=0$. The line of fixed points with arbitrary $\kappa_{0}$ found in lowest order remains valid for $f_{3}=0$, or approximately valid for large $\kappa/k^{2}$. Once in the range of small $k^{2}/\kappa$ the flow of $\kappa$ stops as $k$ decreases towards zero, $\kappa\approx\kappa_{0}+ck^{4}/\kappa_{0}$. The flow equation for $v=k^{2}/\kappa$ has an infrared attractive fixed point for $v=0$. It could have an additional infrared fixed point for finite $v$, for example if $f_{3}$ is negative. In any case, if one starts in the range of attraction of the fixed point at $v=0$ the flow will always reach small values of $v$.
We conclude that the kinetic fixed point (46) is either exact with arbitrary $\kappa_{0}$, or becomes a very good approximation for small enough $k^{2}$. For establishing this result it is crucial that the simplified flow equation and its corrections respect the symmetry of multiplicative scaling of the scalar field. This constitutes an important strength of our approach, which is not easy to realize for other forms of flow equations.
Local gauge symmetry
The simplified flow equation can be seen as a particular version of the gauge invariant flow equation [24, 42]. The central ingredient is the field-dependence of the cutoff function $R_{k}(\varphi)$ in the functional integral (47), (Exact flow equation for field-dependent cutoff). This permits us to choose $R_{k}$ as a function of the covariant Laplacian or other operators that transform homogeneously as tensors with respect to local gauge transformations. In particular, for a gauge invariant flowing action the second functional derivative $\Gamma_{k}^{(2)}$ transforms as a tensor. In consequence, $R_{k}$ should transform as a tensor as well. This requires that $R_{k}$ has to depend on the macroscopic gauge field appearing in covariant Laplacians. The simplified flow equation obeys this requirement for gauge invariance.
We next establish the gauge invariance of the flowing action and the corresponding exact flow equation (61) by using the definition (47), (Exact flow equation for field-dependent cutoff). The difference between the microscopic and the macroscopic gauge field $\chi_{\alpha}-\varphi_{\alpha}$ transforms homogeneously as a vector, and the correlation function $G_{\alpha\beta}$ is therefore a tensor. One concludes that $L_{k,\alpha}$ in eq. (55) transforms as a vector, similar to $\partial\Gamma_{k}/\partial\varphi_{\alpha}$. For a gauge invariant classical action $S[\chi]$ all pieces on the r.h.s. of eq. (Exact flow equation for field-dependent cutoff) are gauge invariant, implying in turn the gauge invariance of the flowing effective action (47) and the exact flow equation (51). The exact correction $Q_{k}$ in eq. (59) transforms as a tensor, implying gauge invariance of the exact flow equation (61). For the expansion of $Q_{k}$ the lowest order (Corrections to the simplified flow equation) transforms as a tensor, leading for the flow equation to gauge invariance of each order in the expansion.
For a continuum formulation the functional integral for local gauge symmetries needs a regularization, typically implemented by a gauge fixing term with associated Faddeev-Popov determinant. We focus on the physical gauge fixing (for Yang-Mills theories this is Landau gauge) which realizes effectively a projection in the physical fluctuation modes [24, 42]. Implementing a cutoff for the gauge-fluctuations and the Faddeev-Popov determinant adds to the flow equations (6), (11), (51), (61), (71) a universal “measure term”. This measure term does not depend on $\Gamma_{k}$, reflecting that it results ultimately from the regularization of the measure in the functional integral. It is a function of covariant derivatives. Its explicit form can be found in ref. [24, 42]. For the case of quantum gravity the coincidence of the measure term with a corresponding term in the background formalism with physical gauge fixing has been discussed explicitly in ref. [37]. The main term in the flow equation besides the measure term involves effectively only the physical fluctuations (e.g. transverse gauge fields for Yang-Mills theories). Correspondingly, the cutoff function involves the projection of the covariant Laplacian on the physical modes, or generalizations thereof as $\Gamma_{k}^{(2)}$.
We conclude that gauge invariance of the flow equation holds whenever $R_{k}$ transforms homogeneously as a tensor. What is new in the present note is the explicit computation of the corrections to the flow equation (11). This correction is of higher loop order. For the case of Yang-Mills theories with a small renormalized gauge coupling $g_{R}$ one expects that the correction is suppressed as compared to the leading term by a factor of $g_{R}^{2}/(16\pi^{2})$. The additional factor of $g_{R}^{2}$ in the beta-function for the running gauge coupling can be seen rather easily. In leading order one has $\Gamma_{k}^{(2)}=Z_{k}D$, with $D$ a (modified) covariant Laplacian and $Z_{k}=g_{R}^{-2}$ the wave function renormalization multiplying in $\Gamma_{k}$ the gauge invariant kinetic term $F_{\mu\nu}F^{\mu\nu}$. With $R_{k}\sim Z_{k}$ the leading term in the flow equation (11) does not involve a factor $Z_{k}$ on the r.h.s., leading to
$\partial_{t}Z_{k}=\frac{c}{16\pi^{2}}\ ,\quad\partial_{t}g_{R}^{2}=-\frac{c}{16\pi^{2}}g_{R}^{4}\ .$ (85)
The correction term $Q_{k}$ in eq (59), (Corrections to the simplified flow equation) contains no factor $Z_{k}$. We may write is as $Q_{k}=\big{[}Z_{k}g_{R}^{2}/(16\pi^{2})\big{]}q_{k}$ which makes the suppression factor $g_{R}^{2}/(16\pi^{2})$ as compared to $\Gamma_{k}^{(2)}+R_{k}$ directly visible. The correction factor contributes to $\partial_{t}g_{R}^{2}$ only in the order $g_{R}^{6}$.
We finally turn to the specific choice of $R_{k}$ leading to the simplified flow equation (6) or its correction (71). If $N(\varphi)$ is chosen to transform homogeneously as a tensor also $\Gamma_{R}^{(2)}$ and $E[\varphi]$ transform as tensors, implying the gauge invariance of the flow equations. In leading order for a Yang-Mills theory one has $N=Z_{k}^{1/2}$ and therefore $E=-cg_{R}^{2}/(16\pi^{2})$. The simple factor $(1-E/2)$ in eq. (6) accounts for more than $90\%$ of the two-loop contribution to $\partial_{t}g_{R}^{2}$ [8]. The combined effects of extended truncations and corrections to eq. (6) are only a rather small effect for small gauge couplings.
Discussion
In this note we propose a version of a simplified functional flow equation which is gauge invariant, admits analytic continuation between euclidean and Minkowski signature, and guarantees that the cutoff removes the critical fluctuations. The first ingredient for this simplified flow equation is the flow equation (11) for a field-dependent cutoff. At this level consistency with all symmetries, in particular local gauge symmetries, can be achieved. We consider eq. (11) here as an approximation, to which we have computed corrections in the form of a loop expansion. This loop expansion should not be confounded with a perturbative expansion of the flow generator in a small coupling. Already the lowest order (11) constitutes a highly non-perturbative setting. The loop expansion only concerns the computations of corrections to the effective cutoff in case where the latter is taken to be field-dependent. It has been argued that eq. (11) may become exact for a further optimization of the flowing effective action [24]. This requires the existence of solutions of rather complex functional differential equations. In view of its complexity we leave this issue open and stick here to the definition (47) (Exact flow equation for field-dependent cutoff) and associated exact flow equation (61).
The second ingredient for the simplified flow equation concerns a particular choice of the cutoff function. This leads to the simple analytic structure and to an effective cutoff for critical fluctuations in a rather general context. This choice of the cutoff function is only possible for a field-dependent cutoff.
Our computation of the form of the correction terms to the flow equation (11) yields further support for the (approximate) validity of computations with gauge invariant flow equations. These flow equations have been used for a rather simple understanding of flowing coupling functions in quantum gravity [27, 37, 38, 39]. We have employed here the simplified flow equation for the computation of the flow of the kinetial in four-dimensional scalar field theories. This has revealed the possible existence of a new kinetic fixed point which realizes the quantum scale symmetry with respect to a multiplicative rescaling of the scalar field at fixed metric and coordinates. A crossover from this fixed point to the trivial fixed point could overcome the triviality problem for scalar quantum field theories in four dimensions.
In our view the main strength of the proposed flow equation resides in its qualitative robustness close to physical intuition, while maintaining a simple form. This makes it suitable for many investigations of qualitative features of non-perturbative systems, while high quantitative accuracy may not be the main goal.
Acknowledgment:
The author would like to thank J. Pawlowski for useful comments.

### v3

@@
 A formulation of the simplified flow equation in terms of renormalized scale invariant fields permits direct access to scaling solutions and associated fixed points.
+Our setting is based on a particular choice of cutoff function which depends on the macroscopic fields.
 Corrections to the simplified flow equation involve a field-dependent modification of the cutoff for which we discuss a systematic expansion.
@@
 We propose an optimization of the definition of the flowing action for which the simplified flow equation can be considered as the starting point of a systematic expansion.
+The key ingredient is the dependence of the cutoff function on the macroscopic fields.
+This induces corrections to the flow equation beyond the one loop form of ref.
+[7].
+We modify the source term in order to cast the corrections into a simple form.
 We compute explicitly the general form of the corrections to the simplified flow equation.
@@
 It includes already all one-loop effects exactly.
-Having a structure similar to the exact flow equation for the effective average action the simplified flow equation also sums up higher loops and can be used for a full non-perturbative analysis.
-The loop expansion only concerns corrections to the precise shape of the infrared cutoff.
+Having a structure similar to the exact flow equation for the effective average action the simplified flow equation also sums up higher perturbative loops and can be used for a full non-perturbative analysis.
+The loop expansion proposed here only concerns corrections to the precise shape of the infrared cutoff.
 Optimization by the choice of a cutoff term which is quadratic in the fluctuation field and independent of the macroscopic field has already a long tradition [23].
@@
 The price to pay is that the one-loop form of the exact flow equation is no longer guaranteed to hold.
-We propose an optimized flow equation which maintains the simple one-loop form.
+We propose an optimized flow equation which maintains formally the simple one-loop form.
 The propagator in the loop involves, however, field-dependent corrections to the effective infrared cutoff which are, in turn, given by a loop expansion.
@@
 For other purposes the effective average action or other forms of optimization may be more appropriate.
-For a single real scalar field $\varphi(x)$ the flowing action $\Gamma_{k}[\varphi]$ is a functional of this “macroscopic field”.
-It flows in dependence on the renormalization scale $k$.
-In the simplest approximation in flat Euclidean space the simplified functional flow equation reads
-$k\partial_{k}\Gamma_{k}[\varphi]=k^{2n}\,\mathrm{tr}\Big{\{}\left(1-\frac{\eta}{2}\right)\left(Z^{-1}\Gamma_{k}^{(2)}[\varphi]+k^{2}\right)^{-n}\Big{\}}\ .$ (1)
-For $d$ even dimensions a typical value is $n=d/2+1$.
-With $\Gamma_{k}^{(2)}$ the matrix of second functional derivatives of $\Gamma_{k}$ the r.h.s.
-of eq.
-(1) involves the $n$-th power of the regularized full field-dependent propagator.
-The quantity $Z$ denotes a suitable wave function renormalization, with anomalous dimension $\eta=-k\partial_{k}\ln Z$.
-Below we will generalize the wave function renormalization $Z$ to a field-dependent quantity.
-This replaces $Z^{-1}\Gamma_{k}^{(2)}$ by a renormalized version of the second functional derivative, which is an important characteristic of our setting.
-Eq.
-(1) has been proposed in [26] for the Wilsonian action which differs conceptually from the flowing effective action.
-In this context it follows from a version of heat kernel regularization.
-Even though heat-kernel evaluations of the simplified flow equation are possible, we do not base our approach on a heat kernel regularization.
-In order to get a first impression of the simplified flow equation one may evaluate the flow of the effective potential according to eq. (1). In momentum space one has
-$\Gamma_{k}^{(2)}(q,q^{\prime})=\frac{\partial^{2}\Gamma_{k}}{\partial\varphi(-q)\partial\varphi(q^{\prime})}\ ,$ (2)
-and $k^{2}$ stands for $k^{2}\delta(q,q^{\prime})$.
-We use conventions with $\int_{q}=\int\text{d}^{d}q(2\pi)^{-d}$, $\delta(q,q^{\prime})=(2\pi)^{d}\delta(q-q^{\prime})$, $\mathrm{tr}A(q,q^{\prime})=\int_{q,q^{\prime}}A(q,q^{\prime})\delta(q^{\prime},q)$, with $\delta^{2}(q,q^{\prime})=\Omega\delta(q,q^{\prime})$ and $\Omega$ the spacetime volume.
-We may take the truncation
-$\Gamma_{k}^{(2)}(q,q^{\prime})=Z\big{(}q^{2}+m^{2}\big{)}\delta(q,q^{\prime})\ ,$ (3)
-with $m^{2}(\varphi)$ given by the second derivative of the scalar potential $m^{2}=\partial^{2}U/\partial\varphi^{2}$. For $d=4$, $n=3$ eq. (1) yields
-$k\partial_{k}\Gamma_{k}=\Omega k^{6}\left(1-\frac{\eta}{2}\right)\int_{q}\big{(}q^{2}+m^{2}+k^{2}\big{)}^{-3}\ .$ (4)
-For a constant scalar field $\varphi$ one has $\Gamma_{k}=\Omega U$, with $U$ the effective potential flowing as
-$k\partial_{k}U=$ $k^{6}\left(1-\frac{\eta}{2}\right)\int_{q}\big{(}q^{2}+m^{2}+k^{2}\big{)}^{-3}$
-$=$ $\frac{\left(1-\frac{\eta}{2}\right)k^{6}}{32\pi^{2}\big{(}m^{2}+k^{2}\big{)}}\ .$ (5)
-Neglecting $\eta$ this yields the same expression as for the standard flow equation with the Litim cutoff [23].
 Simplified flow equation
-We next introduce the general form of the simplified flow equation.
-For several fields $\varphi$ becomes a vector in field space and $\Gamma_{k}^{(2)}$ a corresponding matrix.
 The simplified flow equation is defined by
@@
 $\Gamma_{R}^{(2)}[\varphi]=g^{-1/2}N^{-1}[\varphi]\Gamma_{k}^{(2)}[\varphi]\big{(}N^{T}\big{)}^{-1}[\varphi]\ .$ (2)
-The matrix $N[\varphi]$ generalizes $Z^{1/2}$ to a type of field-dependent wave function renormalization. Correspondingly, the matrix $E[\varphi]$ generalizes $\eta$ in eq. (1). We take
+Here $\varphi$ is the macroscopic field, which is a vector in the space of internal or Lorentz indices, and $\Gamma_{k}^{(2)}[\varphi]$ denotes the second functional derivative of the flowing action with respect to $\varphi$.
+The matrix $N[\varphi]$is a type of field-dependent wave function renormalization.
+Correspondingly, the matrix $E[\varphi]$ is a field-dependent anomalous dimension.
+We take
 $E[\varphi]=-N^{-1}\partial_{t}\big{(}NN^{T}\big{)}\big{(}N^{T}\big{)}^{-1}[\varphi],$ (3)
-such that for a $k$-dependence of $NN^{T}$ of the form $N[\varphi]N^{T}[\varphi]=A[\varphi]k^{-\eta}$ one indeed finds $E[\varphi]=\eta$.
-This includes the field-independent case $NN^{T}=Z$.
+such that for a $k$-dependence of $NN^{T}$ of the form $N[\varphi]N^{T}[\varphi]=A[\varphi]k^{-\eta}$ one finds $E[\varphi]=\eta$.
+The field-dependent factors $N[\varphi]$ become important if the kinetic term in $\Gamma_{k}$ is multiplied by a field-dependent function $\sim NN^{T}$, as for a field-dependent Planck mass in quantum gravity, or the field-dependent “kinetial” for the suggested new fixed point in scalar quantum field theory.
+The trace in eq.
+(1) amounts to a momentum integration or an integration over positions.
+We take $n=3$ for $d=3,4$ and $n=2$ for $d=1,2$, such that the momentum integral is finite for $\Gamma_{k}^{(2)}\sim q^{2}$.
 For an arbitrary metric $g_{\mu\nu}$, with $g=\det(g_{\mu\nu})$, the factor $g^{-1/2}$ takes out the factor $g^{1/2}$ arising from $\Gamma=\int_{x}\sqrt{g}L$.
@@
     The flow equation respects all symmetries, including local gauge symmetries.
-Finally, the flow equation (6) offers an important technical advantage:
+As as technical advantage we may state
   1. (vi)
+    A suitable choice of the field-dependent wave function renormalization $N[\varphi]$ can simplify the form of $\Gamma_{R}^{(2)}$.
+For many examples and a large set of truncations of $\Gamma_{k}$ this can be used to bring $\Gamma_{R}^{(2)}$ to a form involving covariant Laplacians in a simple way.
+One typically aims for settings where $\Gamma_{R}^{(2)}$ is block diagonal and $E$ can be approximated for each block by an anomalous dimension $\eta_{i}$.
+The factors $N^{-1}$ first partly diagonalize $\Gamma_{k}^{(2)}$ and subsequently perform a rescaling.
+Finally, the flow equation (1) offers one more important technical advantage:
+  1. (vii)
     Heat kernel methods can be employed for the evaluation of the flow generator.
@@
 For gauge invariant $\Gamma_{k}$ the second functional derivative is gauge covariant, which makes the gauge invariance of $B(z)$ and therefore the flow generator manifest.
+An interesting approximation to the simplified flow equation considers a simple scalar field $\varphi$ and approximates $NN^{T}$ by a field-independent wave function renormalization $Z$.
+For this approximation in flat Euclidean space the simplified functional flow equation reads
+$k\partial_{k}\Gamma_{k}[\varphi]=k^{2n}\,\mathrm{tr}\Big{\{}\left(1-\frac{\eta}{2}\right)\left(Z^{-1}\Gamma_{k}^{(2)}[\varphi]+k^{2}\right)^{-n}\Big{\}}\ .$ (6)
+This equation has been obtained before from the proper time flow equation [27, 28, 29, 30, 28, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40].
+It has also been proposed in [41] for the Wilsonian action, which differs conceptually from the flowing effective action.
+Our setting is based on a functional integral for the flowing action and we do not use a regularization of heat kernels for the definition of the flowing action.
+The precise connection which leads to the coincidence of eq.
+(6) with a proper time flow equation is not known to us.
+This may constitute an interesting topic of research.
+One also may raise the question if there exist possible connections between the full simplified flow equation (1) and the corrections to it with a regularization of heat kernels.
+The simplified flow equation can also be compared with the background formalism for the effective average action [8].
+The background formalism involves in addition to the macroscopic field $\varphi$ a background field $\bar{\varphi}$ on which the cutoff may depend.
+The flowing action $\Gamma_{k}[\varphi,\bar{\varphi}]$ is then a functional of two fields, for example two gauge fields $A_{\mu}$ and $\bar{A}_{\mu}$.
+Identifying $\bar{A}=A$ the effective action $\Gamma_{k}[A]=\Gamma_{k}[A,A]$ is gauge invariant if the cutoff function involves covariant derivatives with respect to the background field $\bar{A}$.
+The second functional derivative $\Gamma_{k}^{(2)}[A,\bar{A}]$ which appears in the exact one loop form of the flow equation is taken at fixed $\bar{A}$.
+A commonly used approximation replaces $\Gamma_{k}^{(2)}[A,\bar{A}]$ by $\Gamma_{k}^{(2)}[A]$.
+This yields a flow equation with a similar structure as the simplified flow equation.
+In fact, the background formalism equals the flowing action discussed in this note up to a modification of the source term.
+The structure of the corrections to the exact flow equation in the background field formalism is known [8].
+They are rather complicated since two gauge fields $A$ and $\bar{A}$ are involved.
+The corrections to the simplified flow equation have a simpler structure.
 A cutoff depending on the second functional derivative of an effective action and its relation to heat kernels has been discussed for the background field formalism in ref.
 [42, 43].
-For other proposals of non-perturbative flow equations using heat kernels see refs.
-[30, 31, 32, 33, 34, 35, 26, 36].
-As a second technical advantage we may state
-  1. (vii)
-    A suitable choice of the field-dependent wave function renormalization $N[\varphi]$ can simplify the form of $\Gamma_{R}^{(2)}$.
-For many examples and a large set of truncations of $\Gamma_{k}$ this can be used to bring $\Gamma_{R}^{(2)}$ to a form involving covariant Laplacians in a simple way.
-One typically aims for settings where $\Gamma_{R}^{(2)}$ is block diagonal and $E$ can be approximated for each block by an anomalous dimension $\eta_{i}$.
-The factors $N^{-1}$ first partly diagonalize $\Gamma_{k}^{(2)}$ and subsequently perform a rescaling.
+In order to get a first impression of the simplified flow equation one may evaluate the flow of the effective potential according to eq. (6). In momentum space one has
+$\Gamma_{k}^{(2)}(q,q^{\prime})=\frac{\partial^{2}\Gamma_{k}}{\partial\varphi(-q)\partial\varphi(q^{\prime})}\ ,$ (7)
+and $k^{2}$ stands for $k^{2}\delta(q,q^{\prime})$.
+We use conventions with $\int_{q}=\int\text{d}^{d}q(2\pi)^{-d}$, $\delta(q,q^{\prime})=(2\pi)^{d}\delta(q-q^{\prime})$, $\mathrm{tr}A(q,q^{\prime})=\int_{q,q^{\prime}}A(q,q^{\prime})\delta(q^{\prime},q)$, with $\delta^{2}(q,q^{\prime})=\Omega\delta(q,q^{\prime})$ and $\Omega$ the spacetime volume.
+We may take the truncation
+$\Gamma_{k}^{(2)}(q,q^{\prime})=Z\big{(}q^{2}+m^{2}\big{)}\delta(q,q^{\prime})\ ,$ (8)
+with $m^{2}(\varphi)$ given by the second derivative of the scalar potential $m^{2}=\partial^{2}U/\partial\varphi^{2}$. For $d=4$, $n=3$ eq. (6) yields
+$k\partial_{k}\Gamma_{k}=\Omega k^{6}\left(1-\frac{\eta}{2}\right)\int_{q}\big{(}q^{2}+m^{2}+k^{2}\big{)}^{-3}\ .$ (9)
+For a constant scalar field $\varphi$ one has $\Gamma_{k}=\Omega U$, with $U$ the effective potential flowing as
+$k\partial_{k}U=$ $k^{6}\left(1-\frac{\eta}{2}\right)\int_{q}\big{(}q^{2}+m^{2}+k^{2}\big{)}^{-3}$
+$=$ $\frac{\left(1-\frac{\eta}{2}\right)k^{6}}{32\pi^{2}\big{(}m^{2}+k^{2}\big{)}}\ .$ (10)
+Neglecting $\eta$ this yields the same expression as for the standard flow equation with the Litim cutoff [23].
 Relation to exact flow equation
@@
 (11) for field-dependent $B$ and $N$ will be addressed below.
-We may consider eq.
+We may also consider eq.
 (11) as the flow equation in the background formalism for which the correction term discussed in ref.
@@
 We see no obvious reason why similar solutions of the differential equation (18) should not exist for $\gamma\neq 1$.
-This possibly extends to the inclusion of corrections to $\tilde{r}$.
+This possibly extends to the inclusion of corrections to eq.
+(13) or more generally to $r$.
 For an optimal choice of the cutoff function we believe that the dominant corrections to the simplified flow equation arises from the corrections to eq.
@@
 $k\partial_{k}u(\tilde{\rho})=\sum_{i}\frac{2-\eta_{i}}{64\pi^{2}(1+\tilde{m}_{i}^{2})}+(2+\eta)\tilde{\rho}\partial_{\tilde{\rho}}u(\tilde{\rho})-4u\ ,$ (32)
-in accordance with eq. (Simplified functional flow equation).
+in accordance with eq. (Simplified flow equation).
 The sum runs over all particles with renormalized dimensionless mass term $\tilde{m}_{i}^{2}=m_{i}^{2}/k^{2}$.
@@
 Eq.
-(Simplified functional flow equation) holds for an arbitrary constant metric, and therefore both for Euclidean or Minkowski flat space.
+(Simplified flow equation) holds for an arbitrary constant metric, and therefore both for Euclidean or Minkowski flat space.
 The flow of $\tilde{\rho}$-derivatives of $u$ obtains by differentiating both sides of eq.
@@
 $\,+\frac{3c_{2}\partial^{\mu}m^{2}\partial_{\mu}m^{2}k^{6}}{(k^{2}+m^{2})^{4}}\bigg{)}\ .$ (41)
-For homogeneous $m^{2}$ we recover the flow of the potential in eq. (Simplified functional flow equation).
+For homogeneous $m^{2}$ we recover the flow of the potential in eq. (Simplified flow equation).
 With $m^{2}=\partial_{\varphi}^{2}U/K-C$ one employs for the second order of the derivative expansion